Use the Pythagorean theorem to find the unknown lengths.
The sides of a square are 36inches. Find the diagonal of the square. Round your answer to 2 decimal places.

Answer by rfer(16322)

36^2+36^2=c^2
1296+1296=c^2
2592=c^2
sqrt 2592=50.91 in